A 40.0-kg student climbs up a ladder to the roof of a building. The gravitational potential energy of the student at this point is 2.00 x 103 J. How high is the student above the ground?


1
Expert's answer
2021-05-30T13:31:59-0400

By definition, the potential energy is:


"E = mgh"

where "m = 40kg" is the mass of the student, "g = 9.81m\/s^2" is the gravitational acceleration, "h" is the height. Substituting "E = 2\times 10^{3}J" and expressing "h", obtain:


"h = \dfrac{E}{mg} = \dfrac{2\times 10^3}{40\cdot 9.81} \approx 5.10m"

Answer. 5.10 m.